OpenStudy (anonymous):

Q and R are independent events. If P(Q)=0.8 and P(R)=0.2 find P(Q and R). (a). 0.16 (b). 0.84 (c). 1

OpenStudy (kropot72):

An important property of independent events is that the probability of both occurring is the product of their probabilities. This is written as the formula: \[P(A\cap B)=P(A) \times P(B)\]

OpenStudy (anonymous):

I am still confused im not sure what you mean by that

OpenStudy (anonymous):

For instance where did you get the A and the B? Can you explain and break it down please

OpenStudy (kropot72):

So for your question: \[P(Q \cap R)=P(Q) \times P(R)\ ..........(1)\] Now you just plug the given values into (1)

OpenStudy (anonymous):

so so i multiply the given values 0.8 and 0.2?

OpenStudy (kropot72):

Correct!

OpenStudy (anonymous):

).16? Thank you SOOOO MUCH!

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

:) I have a quiz im taking for summer school and theres no teacher so this means a lot

OpenStudy (kropot72):

Glad to be able to help.